OpenStudy (anonymous):

-6x+2=-4(2x+4) can someone do the stps

OpenStudy (saifoo.khan):

-6x+2=-4(2x+4) -6x + 2 = -8x - 16 Bring variables on one side and numbers on other, -6x + 8x = - 16 -2 2x = -18 Now divide 2 from both sides, x = -18/2 x = -9

OpenStudy (saifoo.khan):

welks.